OpenStudy (anonymous):

Suppose f(π/3) = 3 and f '(π/3) = −7, and let g(x) = f(x) sin x and h(x) = (cos x)/f(x). Find the h'(x)

OpenStudy (anonymous):

i know its the quotient rule this is what i got: -7cos(x)+sin(x)-3/(cosx^2)

OpenStudy (anonymous):

but thats incorrect

OpenStudy (anonymous):

you're sure it's not find h'(π/3) ?

OpenStudy (anonymous):

yeah its h'(pi/3) sorry

OpenStudy (anonymous):

why is it incorrect? did i make a simple error?

OpenStudy (anonymous):

Let's give this a go: h(x) = cos(x) / f(x) derivative (recall the quotient rule) h'(x) = [ f(x) * (-sin(x)) - cos(x)*f'(x) ] / [ f(x) ]^2 simplify h'(x) = [ -sin(x)*f(x) - cox(x)*f '(x) ] / [ f(x) ]^2 h'(π/3) = [ -sin(π/3)*f(π/3) - cox(π/3)*f '(π/3) ] / [ f(π/3) ]^2 h'(π/3) = \[ -(\sqrt{3}/2)*(3) -(1/2)*(-7) /(3)^2\] h'(π/3) = \[(-3\sqrt{3}/2+7/2)/9\] And you can further simplify if you want, I'll stop there.

OpenStudy (anonymous):

thank you